# Simulations of Classical Three-Body Thermalization in One Dimension

### v1

@@ II Gas and Thermalization Model
where $d$ is the pairwise separation, $l_{0}$ is the basic unit of length at which the potential energy is equal to some interaction strength $\epsilon>0$, and $n$ is an even integer $\geq 2$.
The locations of the particles are parameterized by the angles $\phi_{i}$ and the mutual distance between particles $i$ and $j$ is
@@
where $v_{th}=\sqrt{kT/m}$ is the velocity scale set by the temperature. $k$ is the Boltzmann constant.
In this work, we assume high enough $T$ and small enough number density, $N/(2\pi R)$, so that the average kinetic energy is much greater than the potential energy (unlike in Ref. [5], where the energies are comparable). In this regime, correlations between positions of the particles, higher order collisions, and collective effects can be ignored as in Ref. [16].
@@
We choose a ring to represent the 1D system instead of a line so that particles do not escape to infinity under repulsive forces. Modeling a line would require a confining potential which leads to the particles at the edges experiencing a different environment from those in the center of the range. We are aware of the two artifacts of the ring: curvature and periodic effects. To simulate a particular linear number density, $\rho$, we tackle the ring effects by doubling $R$ and the number of thermal particles, $N-1$, until our results converge.
We consider a perturbation to the thermal state by starting the $N^{th}$ particle at velocity $v^{\prime}$ as in Fig. 1. We refer to such a particle as the ‘special’ particle and its velocity as the ‘special velocity’, while we refer to the rest of the particles as ‘thermal’. The velocity probability density function of the total system is
@@
Our goal is to study the evolution of $h$. According to the Boltzmann $H$-theorem [10], the steady state distribution $f(v,\infty)$, and therefore $h(v,\infty)$, is the equilibrium distribution $f_{0}$. This is strictly true in the thermodynamic limit $N\rightarrow\infty$. For finite $N$, however, the steady state has a slightly different temperature $T_{f}\simeq(N-1+u^{\prime 2})T/N\approx T$ for large $N$, which is a result of energy conservation. We run many MD simulations where the $N-1$ thermal particles are randomly chosen from a thermal distribution. Averaging over these many runs gives an average $h(u,t)$.
From our MD simulations (as described in Appendix A), the average $h(u,t)$ starting from $\delta(u)$ is shown for two times in Fig. 2. As expected, the perturbation’s peak height decreases and its width increases as the system evolves. The times shown are early in the thermalization process where the population only spreads to small velocities ($|u|<0.06$). However, the cusp feature of the delta function is maintained throughout this time period, which is an indication that the gas is far from equilibrium. The parameters of the simulated gas are $l_{0}=1\,\mu m$, $\epsilon=1\,eV$, $kT=100\,meV$, $m=m_{e}$ (mass of the electron), $n=6$, $N=41$, and $R=6400\,\mu m$. The times shown are $t_{f}=960,3200\,ns=0.48,1.6\times 10^{-4}/\Gamma$ ($\Gamma$ being the rate of thermalization defined in Eq. (8)).
Figure 2: MD simulated evolution of the perturbation initially $\delta(u)$ for parameters $N=41$ and $R=6400\,\mu m$. The times shown are $t_{f}=960,3200\,ns=0.48,1.6\times 10^{-4}/\Gamma$, and the $x$-axis is the scaled velocity $u$. Comparing the distribution at these two time instants and at $t=0$ (not shown), the perturbation width increases with time as its height drops.
@@
This definition of $\Gamma$ implies that if the perturbation continues spreading at the initial rate, then it takes time of order $1/\Gamma$ for the perturbation $h(u)$ to reach variance $\sim 1$, i.e., the special particle reaching the temperature of the bath. In the rest of the paper, we show how $\Gamma$ depends on the system parameters $\{R,N,m,T,l_{0},\epsilon,n,u^{\prime}\}$. To extract this dependence, we do MD simulations for different parameter sets; however, as argued in Appendix A, such $N$-body simulations are computationally very expensive when $N$ is large or the propagation time, $t_{f}$, is large.
@@ III Boltzmann Evolution of the Gas
$h_{i}(t+dt)\approx h_{i}(t)+dt\sum_{j=-\infty}^{j=\infty}K_{v_{j}\rightarrow v_{i}}h_{j}(t),$ (11)
where $h_{i}\Delta v$ is the population in bin $v_{i}$,
To ensure that the population (particle number) is conserved, the scattering rate out of velocity bin $v_{i}$ ($K_{v_{i}\rightarrow v_{i}}$) must equal the scattering rate into any other velocity $v_{j}$. That is,
$K_{v_{i}\rightarrow v_{i}}=-\sum_{j\neq i}K_{v_{i}\rightarrow v_{j}}.$ (12)
where the minus sign indicates that population is lost from bin $v_{i}$.
@@ IV.1 Scaling Behavior of the $N$-body Gas
From the definition of $l$ in Eq. (15), $kT$ scales as $l^{-n}$. Furthermore, the density $\rho$ of the particles for a given $N$ scales as $R^{-1}$. Thus, $c$ scales as $T^{-1}\rho^{n}$. If one scales ${\rho}{\rightarrow}a\rho$ and $T{\rightarrow}a^{n}T$, $c$ remains the same. Therefore, the scaled dynamics (Eq. (IV.1) alongside the scaled initial conditions) remains the same . Given this constraint, one can show that the most general time scale must be proportional to $\rho^{-s}T^{-(\frac{1}{2}-\frac{s-1}{n})}$, where $s$ is an integer. These time scales are also obtainable from dimensional analysis of the system’s parameters. The most general rate $\gamma$ is thus governed by terms of the form (rewritten using $l$ and $v_{th}$)
@@
Cases where $s\geq 0$ have physical interpretations. $s=0$ gives a time scale $t_{c}={l}/v_{th}$ which is proportional to the interaction time $during$ a binary collision. $s=1$ gives the dynamical time scale $t_{d}$ which is proportional to the mean time $between$ binary collisions. Similarly, $s=2$ gives the rate of ternary interactions. The probability that two particles existing in an interaction region of length $l$ scales as ${\rho}l$, and the rate of a third particle entering this region to interact with the other two particles is ${\rho}v_{th}$. Thus the ternary interaction rate scales as $\rho^{2}lv_{th}$.
@@ IV.2 Scaling of the Collision Kernel $K$
$\Gamma_{coin}$ $=$ $\frac{probability\,of\,two\,arrivals}{t_{ob}}$ (22)
@@
Figure 4: $G_{0\rightarrow u}$ for different values of $T$ and $\rho l$. The height and shape of all curves are identical which demonstrates the universal scaling. A fitting function $\frac{\alpha}{|u|^{\beta}}e^{-u^{2}/2\sigma^{2}}$ is also plotted and explained in Sec. IV.5.
@@ IV.3 MD and Boltzmann Evolution Comparison
The spread of the perturbation in Fig. 6 indicates that its variance (which is proportional to the energy) is growing with time which is shown in Fig. 7. The linear evolution of the perturbation’s energy in the time range considered suggests modelling the thermalization process as a random walk of the velocity $u^{\prime}$ of the special particle. The standard deviation of $h$ is proportional to $\sqrt{t}$, just like the standard deviation of displacement in a random walk is proportional to the square root of the number of steps taken.
@@ IV.4 G for a moving special particle $u^{\prime}\neq 0$
while the ratio between the Maxwell-Boltzmann population at the corresponding bins is ${e^{-{2.96^{2}}/{2}}}/{e^{-{3.00^{2}}/{2}}}=1.13.$
@@ IV.5 Dependence of $G$ on the Potential Power $n$
as shown in Fig. 4 for $n=6$. The parameters $\alpha$, $\beta$, and $\sigma$ are shown for several values of $n$ in Tab.1. The $\alpha$ parameter is proportional to the over all scattering rate $K_{0\rightarrow 0}$, Eq. (12), and the power $\beta$ measures how fast the kernel drops with increasing $|u|$ near $u=0$ as in Fig. 4. We observe that both $\alpha$ and $\beta$ decrease as $n$ increases (i.e., when the potential is becoming steeper and approaching hard walls).
@@
Figure 11: Scaled kick $u_{1f}$ as a function of the scaled delay $\Delta t$ (in units of $l/v_{th}$) between launches. The symmetric case is for $u_{2i}=1$ and $u_{3i}=-1$, while the asymmetric case is for $u_{2i}=1.1$ and $u_{3i}=-0.9$. The curve for the asymmetric case was shifted horizontally for clarity. Both curves have long tails and peak at $u_{1f,max}$ $\sim 0.03$.
@@ Appendix A Molecular Dynamics Simulation
In this appendix, we provide details and analysis of the MD simulations. To simulate the gas of $N$ particles in Fig. 1, the initial velocities are chosen as per Eq. (5), whereas the initial locations are chosen randomly according to a relative potential energy Boltzmann factor. We evolve the system from time $t_{i}=0$ to time $t_{f}$ according to Eq. (IV.1) using Runge–Kutta methods with adaptive time step [21]. At $t_{f}$, we subtract the background thermal distribution as per Eq. (5) to single out the perturbation $h(v,t_{f})$. We average over many trajectories and a window of time around $t_{f}$ to reduce the statistical noise.
The number of terms in the force calculation in Eq. (IV.1) scales like $N^{2}$ for a single time step, which makes it difficult to simulate more than 30 particles. To decrease the computational time, we utilize that the inverse power potential in Eq. (1) is relatively ‘short-range’ for $n\geq 2$ and small density $\rho$. In this regime, particles are only interacting significantly with their close neighbors, while the interaction with further particles can be ignored. This allows us to do nearest neighbors calculations for the force, which results in a time complexity that scales like $N$. The neighbors are selected by ordering the particles according to their initial locations. For 3 nearest neighbors, for example, particle number 5 experiences a force from particles numbered 2-4 and 6-8. Since the potential is infinitely repulsive at short distance, the particles cannot pass through each other and the particle order is fixed. We checked that the results converged for 3 nearest neighbors which was used for all MD simulations in this paper. Using this method, we could simulate more than 160 particles.

### v2

@@ II Gas and Thermalization Model
where $d$ is the pairwise separation (the 2D distance in the plane of the ring), $l_{0}$ is the basic unit of length at which the potential energy is equal to some interaction strength $\epsilon>0$, and $n$ is an even integer $\geq 2$.
The locations of the particles are parameterized by the angles $\phi_{i}$ and the 2D distance (the length of the chord) between particles $i$ and $j$ is
@@
We choose a ring to represent the 1D system instead of a line so that particles do not escape to infinity under repulsive forces. Modeling a line would require a confining potential which leads to the particles at the edges experiencing a different environment from those in the center of the range. Although the aim of the paper is studying the gas in 1D, we use the 2D distance along the chord in Eq. (2) instead of the 1D distance along the arc so that the torque would be continuous in $\phi$ and would be 00 for particles on opposite sides of the ring. It also has the added benefit of smoothly cutting off the interaction between two particles as $|\phi_{i}-\phi_{j}|$ increases. This introduces curvature effects which distort from a true 1D system. However, these effects become negligible at large $R$. The choice of a finite ring with periodic boundary also introduces finite size effects that diminish at large $R$ and $N$. To simulate a particular linear number density, $\rho=N/R$, we decrease the ring effects by doubling $R$ and $N$ until our results converge.
@@
where $v_{th}=\sqrt{kT/m}$ is the velocity scale set by the temperature. $k$ is the Boltzmann constant. The initial locations are chosen randomly according to a relative potential energy Boltzmann factor $exp(-PE/kT)$, where $PE$ is the total potential energy (second term in Eq. (3)). In this way, a configuration with a very close group of particles occurs rarely and, therefore, rejected with a high probability in our simulations. Also, the number density $\rho$ is homogeneous on average. In our MD simulations, we ensure uniform density by averaging the results over many runs with different initial conditions. We do not consider effects of density variations along the ring (which lead to the diffusion of the particles) in this paper. However, we will show how the relaxation rate depends on the overall density.
In this work, we assume high enough $T$ and small enough number density, $\rho$, so that the average kinetic energy is much greater than the potential energy (unlike in Ref. [5], where the energies are comparable). In this regime, correlations between positions of the particles can be ignored. At the other extreme of low $T$ and high $\rho$, position correlations become important because the system condenses to a solid with periodic arrangement of particles.
@@
We consider a perturbation to the thermal state by forcing one of the particles (e.g. the $N^{th}$ particle) to start at a particular velocity $v^{\prime}$ as in Fig. 1. While all the rest of the phase variables obey the thermal distribution statistics, this particle’s velocity will disobey the Maxwell Boltzmann distribution in Eq. (4) and will rather have a delta function distribution centered at $v^{\prime}$. We refer to such a particle as the ‘special’ particle and its velocity as the ‘special velocity’, while we refer to the rest of the particles as ‘thermal’. The velocity probability density function of the total system is
@@
Our goal is to study the evolution of $h$. According to the Boltzmann $H$-theorem [10], the steady state distribution $f(v,\infty)$, and therefore $h(v,\infty)$, is the equilibrium distribution $f_{0}$. This is strictly true in the thermodynamic limit $N\rightarrow\infty$. For finite $N$, however, the steady state has a slightly different temperature $T_{f}\simeq(N-1+u^{\prime 2})T/N\approx T$ for large $N$, which is a result of energy conservation.
@@
This definition of $\Gamma$ implies that if the perturbation continues spreading at the initial rate, then it takes time of order $1/\Gamma$ for the perturbation $h(u)$ to reach variance $\sim 1$, i.e., the special particle reaching the temperature of the bath. This definition is introduced to enable the quantification of the notion of the thermalization rate.
We choose this particular type of perturbation because it is easy to track and characterize during the simulations. The initial perturbed distribution is a Gaussian with a large narrow spike at $v^{\prime}$. When the number of particles is large, the Gaussian background is mostly unaffected during the evolution. On the other hand, during the early stages of the thermalization the spike spreads out but remains significantly higher than the background Gaussian. It is therefore possible, during early times, to track the evolved perturbation by only tracking a small range of velocities (we choose a range of $0.2v_{th}$ centered around $v^{\prime}$) and subtracting the background to get $h$. This eliminates the need to make a histogram of all the velocities and makes it possible to get a clearer evolution of the perturbation. Although our choice is mostly for ease of computations, it will be shown below that the relaxation rate, $\Gamma$, is only different by a small numerical factor for different types perturbations.
We run many MD simulations (as described in detail in Appendix A) where the $N-1$ thermal particles are randomly chosen from a thermal distribution. Averaging over these many runs gives an average $h(u,t)$. From our simulations, the average $h(u,t)$ starting from $\delta(u)$ is shown for two times in Fig. 2. As expected, the perturbation’s peak height decreases and its width increases as the system evolves. The times shown are early in the thermalization process where the population only spreads to small velocities ($|u|<0.06$). However, the cusp feature of the delta function is maintained throughout this time period, which is an indication that the gas is far from equilibrium. The parameters of the simulated gas are $l_{0}=1\,\mu m$, $\epsilon=1\,eV$, $kT=100\,meV$, $m=m_{e}$ (mass of the electron), $n=6$, $N=41$, and $R=6400\,\mu m$. The times shown are $t_{f}=960,3200\,ns=0.48,1.6\times 10^{-4}/\Gamma$, where $\Gamma$ is computed in Eq. (23).
Figure 2: MD simulated evolution of the probability distribution of the perturbation $h(v,t)$ which is initially localized at $v=0$. The two times shown are $0.48$ and $1.6$ (×$(\times 10^{-4}/\Gamma$) which are very early in the thermalization process. The $x$-axis is the scaled velocity $u=v/v_{th}$. Comparing the distribution at these two time instants and at $t=0$ (a delta function of $u$), the perturbation width increases with time as its height drops. The fluctuations at the edges of the curves are due to statistical noise.
In the rest of the paper, we show how $\Gamma$ depends on the system parameters $\{R,N,m,T,l_{0},\epsilon,n,u^{\prime}\}$. To extract this dependence, we do MD simulations for different parameter sets; however, as argued in Appendix A, such $N$-body simulations are computationally very expensive when $N$ is large or the propagation time, $t_{f}$, is large.
@@ III Boltzmann Evolution of the Gas
$h_{i}(t+dt)\approx h_{i}(t)+dt\sum_{j=-\infty,j\neq i}^{j=\infty}\Delta v\left[K_{v_{j}\rightarrow v_{i}}h_{j}(t)-K_{v_{i}\rightarrow v_{j}}h_{i}(t)\right],$ (11)
where $h_{i}\Delta v$ is the population in bin $v_{i}$. The positive term in the square brackets represents the flow of population to bin $v_{i}$ from all other bins, while the negative term represents the flow out of bin $v_{i}$ to all other bins.
@@ III.1 Monte Carlo Simulation of $K$
This idea of tracking a particle in contact with a bath in thermal equilibrium is similar to the analysis of Brownian motion [18]. But instead of following the position of a tagged particle to see how it diffuses in space (like in a random walk), we track the special velocity and see how it changes due to 3-body collisions. By studying the velocity fluctuations and their rates, we can infer the relaxation rates of macroscopic perturbations. This is similar to using the time correlation function of a particle velocity in Brownian motion to calculate the diffusion coefficient[18]. Deducing the thermalization rate from the velocity changes is an application to the Onsager regression hypothesis which relates macroscopic non-equilibrium disturbances to microscopic fluctuations in the corresponding equilibrium system[19].
@@ IV.1 Scaling Behavior of the $N$-body Gas
From the definition of $l$ in Eq. (14), $kT$ scales as $l^{-n}$. Furthermore, the density $\rho$ of the particles for a given $N$ scales as $R^{-1}$. Thus, $c$ scales as $T^{-1}\rho^{n}$. If one scales ${\rho}{\rightarrow}a\rho$ and $T{\rightarrow}a^{n}T$, $c$ remains the same. Therefore, the scaled dynamics (Eq. (IV.1) alongside the scaled initial conditions) remains the same . Given this constraint, one can show that any time unit of the form $\rho^{-s}T^{-(\frac{1}{2}-\frac{s-1}{n})}$ would be invariant under our scaling symmetry. An invariant rate unit $\gamma$ is thus governed by terms of the form (rewritten using $l$ and $v_{th}$)
@@
Physically relevant time scales could be identified with integer values of $s\geq 0$. $s=0$ gives a time scale $t_{c}={l}/v_{th}$ which is proportional to the interaction time $during$ a binary collision. $s=1$ gives the dynamical time scale $t_{d}$ which is proportional to the mean time $between$ binary collisions. Similarly, $s=2$ gives the rate of ternary interactions. The probability that two particles existing in an interaction region of length $l$ scales as ${\rho}l$, and the rate of a third particle entering this region to interact with the other two particles is ${\rho}v_{th}$. Thus the ternary interaction rate scales as $\rho^{2}lv_{th}$.
@@ IV.2 Scaling of the Collision Kernel $K$
$\Gamma_{coin}$ $=$ $\frac{\mathrm{probability\,of\,two\,arrivals}}{t_{ob}}$ (21)
@@
Figure 4: The dimensionless scaled collision kernel $G_{0\rightarrow u}$ for different values of $T$ and $\rho l$. The height and shape of all curves are identical which demonstrates the universal scaling. A fitting function $\frac{\alpha}{|u|^{\beta}}e^{-u^{2}/2\sigma^{2}}$ is also plotted and explained in Sec. IV.5.
@@ IV.3 MD and Boltzmann Evolution Comparison
The spread of the perturbation in Fig. 6 indicates that its variance (which is proportional to the energy) is growing with time which is shown in Fig. 7. The linear evolution of the perturbation’s energy in the time range considered indicates the thermalization rate defined through Eq. (8) is approximately constant with time and does not depend strongly on the shape of perturbation. It also suggests modelling the thermalization process as a random walk of the velocity $u^{\prime}$ of the special particle. The standard deviation of $h$ is proportional to $\sqrt{t}$, just like the standard deviation of displacement in a random walk is proportional to the square root of the number of steps taken.
@@ IV.4 G for a moving special particle $u^{\prime}\neq 0$
while the ratio between the Maxwell-Boltzmann populations at the corresponding bins is ${e^{-{2.96^{2}}/{2}}}/{e^{-{3.00^{2}}/{2}}}=1.13.$
@@ IV.5 Dependence of $G$ on the Potential Power $n$
as shown in Fig. 4 for $n=6$. The parameters $\alpha$, $\beta$, and $\sigma$ are shown for several values of $n$ in Tab.1. The $\alpha$ parameter is proportional to the over all scattering rate, and the power $\beta$ measures how fast the kernel drops with increasing $|u|$ near $u=0$ as in Fig. 4. We observe that both $\alpha$ and $\beta$ decrease as $n$ increases (i.e., when the potential is becoming steeper and approaching hard walls).
@@
Figure 11: Scaled kick $u_{1f}$ as a function of the scaled delay ($\Delta tv_{th}/l$) between launches. The symmetric case is for $u_{2i}=1$ and $u_{3i}=-1$, while the asymmetric case is for $u_{2i}=1.1$ and $u_{3i}=-0.9$. The curve for the asymmetric case was shifted horizontally for clarity. Both curves have long tails and peak at $u_{1f,max}$ $\sim 0.03$.
@@ V Conclusion
Data for the figures used in this publication are available from the Purdue University Research Repository [23].
@@ Appendix A Molecular Dynamics Simulation
In this appendix, we provide details and analysis of the MD simulations. To simulate the gas of $N$ particles in Fig. 1, the initial velocities are chosen as per Eq. (5), whereas the initial locations are chosen randomly according to a relative potential energy Boltzmann factor. We evolve the system from time $t_{i}=0$ to time $t_{f}$ according to Eq. (IV.1) using Runge–Kutta methods with adaptive time step [24]. At $t_{f}$, we subtract the background thermal distribution as per Eq. (5) to single out the perturbation $h(v,t_{f})$. We average over many trajectories with different initial conditions and average over a window of time around $t_{f}$ to reduce the statistical noise.
The number of terms in the force calculation in Eq. (IV.1) scales like $N^{2}$ for a single time step, which makes it difficult to simulate more than 30 particles. To decrease the computational time, we utilize that the inverse power potential in Eq. (1) is relatively ‘short-range’ for $n\geq 2$ and small density $\rho$. In this regime, particles are only interacting significantly with their close neighbors, while the interaction with further particles can be ignored. This allows us to do nearest neighbors calculations for the force, which results in a time complexity that scales like $N$. This approximation is similar to truncating the potential at certain particle separation as usually used in the MD simulation for the Lennard-Jones potentials[25]. The neighbors are selected by ordering the particles according to their initial locations. For 3 nearest neighbors, for example, particle number 5 experiences a force from particles numbered 2-4 and 6-8. Since the potential is infinitely repulsive at short distance, the particles cannot pass through each other and the particle order is fixed. We checked that the results converged for 3 nearest neighbors which was used for all MD simulations in this paper. Using this method, we could simulate more than 160 particles.